# Fiducial Inference for Random-Effects Calibration Models: Advancing Reliable Quantification in Environmental Analytical Chemistry

### v2

@@ 1 Introduction
A critically important problem in interlaboratory calibration is the accurate interval estimation of the unknown true concentration of an analyte in the context of quantitative analysis in analytical chemistry. The problem is especially challenging when dealing with heterocedastic measurement errors at different concentration levels of the analyte. In practice, analytical measurements are routinely used for various purposes, often without considering the associated level of uncertainty. For instance, in environmental monitoring programs, the concentration of low-level environmental pollutants in a single sample is directly compared with regulatory standards without accounting for the uncertainty in those measurements (see Gibbons, 1994, for reviews of statistical issues in environmental monitoring applications). The uncertainty can be quantified by providing confidence bounds for the true concentration level, given measured concentrations obtained under a hierarchical structure, such as concentrations measured across different laboratories. It is common in environmental monitoring to use multiple laboratories for routine monitoring or to cross-validate an anomalous result. This nested structure, where the laboratory is considered a random factor, complicates the computation of confidence bounds for the calibration problem.
@@
Previous work in this area for linear models includes studies by Osborne (1991); Eberhardt and Mee (1994); Johnson and Krishnamoorthy (1996); Cameron et al. (2022). In addition, data analysis using the Rocke-Lorenzato model and its derivatives typically employs likelihood-based methods, maximizing the integrated likelihood to estimate parameters and their asymptotic variances. Bhaumik and Gibbons (2005) obtained moment-based estimators of the model parameters, studied their properties, and used them to construct confidence intervals using data from multiple laboratories. However, conventional large-sample confidence intervals derived from these methods often exhibit inadequate coverage probabilities, unless sample sizes are large. In response to this limitation, Zhao et al. (2021) proposed a modified likelihood ratio approach using a parametric bootstrap based on a standardized signed log-likelihood ratio test to improve confidence interval accuracy. In spite its improved performance, the coverage probabilities were still unsatisfactory for some parametric choices when multiple laboratories are involved. We shall thus explore a fiducial approach under the Rocke-Lorenzato model and its generalizations based on measurements made by a single laboratory or split samples from multiple laboratories. Based on fairly extensive simulations, we shall demonstrate that the fiducial approach outperforms all other available approaches, including the modified likelihood ratio procedure due to Zhao et al. (2021) and bootstrap-based approaches, in terms of meeting the coverage probability requirement, and yet providing confidence intervals having comparable or better expected widths. We shall also see that the proposed fiducial approach not only has better accuracy, but is also computationally much more efficient than some of the competing methods, especially the methodology developed in Zhao et al. (2021). The fiducial approach has actually seen renewed interest in recent years; Hannig et al. (2016) and Bhaumik et al. (2021). Earlier, Weerahandi (1993), Xu and Wang (2008), Krishnamoorthy and Mathew (2003) and Bhaumik et al. (2013) developed and successfully applied the fiducial approach, referring to a generalized pivotal quantity. We also want to point out that the Rocke-Lorenzato model and its derivatives have random effects that enter the model in a non-linear fashion, and we have been able to develop the fiducial approach in this non-linear scenario. We have also explored the performance of the point estimate by fiducial mode.
@@ 2 Motivating Examples
The goals of the above projects were (a) to estimate the calibration curve for each laboratory, (b) estimate the true unknown analyte concentration, and construct corresponding 100$(1-\alpha)$confidence intervals based on data from multiple laboratories, (c) illustrate the use of the confidence interval in environmental monitoring applications where measurements are routinely compared to regulatory thresholds (RT) without incorporating uncertainty in their estimated concentration.
@@ 4 Fiducial Approach
The notion of fiducial inference is based on a fiducial quantity, introduced by Fisher in the 1930s. Unlike Bayesian inference, it provides probabilistic statements about the unknown parameters without relying on prior distributions. The fiducial quantity is a function of random variables and corresponding observed values such that (i) the distribution of the fiducial quantity given the data does not depend on any unknown parameters, and (ii) the observed value of the fiducial quantity obtained by replacing each random variable by the corresponding observed data is free of the nuisance parameters, and is often equal to the parameter of interest. Applications based on fiducial quantities are available in numerous articles. In particular, in the context of lognormal distributions, such applications are given in Krishnamoorthy and Mathew (2003), Bebu and Mathew (2008), and Bhaumik et al. (2021).
@@
Remark: Interest in the fiducial approach was revived by the seminal works of Tsui and Weerahandi (1989) and Weerahandi (1993); a fiducial quantity was referred to as a generalized pivotal quantity in the latter work. Several later authors have developed and successfully applied the fiducial approach, referring to the fiducial quantity as a generalized pivotal quantity; see, for example, the articles by Krishnamoorthy and Mathew (2003), Bebu and Mathew (2008), Xu and Wang (2008), Bhaumik et al. (2013) and the book by Krishnamoorthy and Mathew (2009). The fiducial quantities we have considered in our work are in the spirit of the generalized pivotal quantities considered by these authors. We have referred to them as fiducial quantities. However, what we have developed are approximate fiducial quantities since their implementation relies on the consistency of the maximum likelihood estimate of the model parameter $\sigma_{\eta}$. In what follows, we derive fiducial quantities for each model parameter and for an unknown true concentration.
@@ 4.1.1 Initial Fiducial Quantities for $\alpha_{i}$ and $\sigma_{\epsilon}$
First, we obtain the fiducial quantities of $\alpha_{i}$, $i=1,...,q$, and $\sigma_{\epsilon}$ based on the data corresponding to $x_{j}=0$. In the next subsections, we update these fiducial quantities based on the complete data (corresponding to all concentrations $x_{j}$) to obtain updated fiducial quantities of $\alpha_{i}$, $i=1,...,q$, and $\sigma_{\epsilon}$. We assume $x_{1}=0$ and for $j=1$ the model (2) can be written as $y_{i1k}=\alpha_{i}+\sigma_{\epsilon}z_{i1k}^{\epsilon}.$ By taking an average over all replications from laboratory $i$ we obtain, $\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}=\alpha_{i}+\sigma_{\epsilon}\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}.$ Thus, we obtain, $\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}=\sigma^{2}_{\epsilon}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}.$ By summing both sides over all laboratories, so as to utilize information across all labs, we get $\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}=\sigma^{2}_{\epsilon}\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}$. Thus, an initial fiducial quantity for $\sigma_{\epsilon}$, say $\tilde{\sigma}^{init}_{\epsilon}$, can be obtained by solving the above equation for $\sigma_{\epsilon}$: $\tilde{\sigma}^{init}_{\epsilon}=\sqrt{\frac{\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}}{\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}}}.$ In the expression for $\tilde{\sigma}^{init}_{\epsilon}$ given above, the $y_{i1k}$s are to be treated as fixed, and the random quantities are the random variable $z_{i1k}^{\epsilon}$. An initial fiducial quantity for $\alpha_{i}$ can be obtained as $\tilde{\alpha}^{init}_{i}=\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}-\tilde{\sigma}^{init}_{\epsilon}\times\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}.$
Remark. In the expression for $\tilde{\sigma}^{init}_{\epsilon}$ given above, the quantity $\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}$ clearly has a chisquare distribution. However, while implementing the fiducial methodology, we will be generating the independent standard normal quantities $z_{i1k}^{\epsilon}$. The reason is that these quantities also appear in later expressions; see for example equation (3).
@@ 4.1.2 Fiducial Quantities for $\beta_{i}$
We construct a total of $q$ structural equations, one for each $\beta_{i}$, $i=1,...,q$. These equations are linear in each $\beta_{i}$, $i=1,...,q$; thus, solutions are obtained in closed forms. We use (2) to construct structural equations for each $\beta_{i}$, $i=1,...,q$, by averaging (2) over all replications corresponding to different concentrations, $\frac{1}{r}\sum\limits_{j=1}^{r}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}y_{ijk}=\alpha_{i}+\beta_{i}\frac{1}{r}\sum\limits_{j=1}^{r}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}x_{j}\exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\frac{1}{r}\sum\limits_{j=1}^{r}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}z_{ijk}^{\epsilon},\ i=1,...,q.$ Parameters $\alpha_{i}$, $i=1,...,q$, and $\sigma^{2}_{\epsilon}$ in these structural equations are replaced by their initial fiducial quantities, $\tilde{\alpha}^{init}_{i}$, $i=1,...,q$, and $\tilde{\sigma}^{init}_{\epsilon}$ respectively. We replace $\sigma^{2}_{\eta}$ with its consistent point estimate $\hat{\sigma}^{2}_{\eta}$ which can be obtained by the maximum likelihood or the method of moments proposed by Bhaumik et al. (2004). Using $\hat{\sigma}^{2}_{\eta}$, the fiducial quantities that we exhibit for the $\beta_{i}$, $i=1,...,q$, are only approximate fiducial quantities; however, we shall henceforth refer to the approximate fiducial quantities simply as fiducial quantities. Thus, the fiducial pivots for $\beta_{i}$, $i=1,...,q$, can be obtained as
$\tilde{\beta}_{i}=\frac{\sum\limits_{j=1}^{r}\sum\limits_{k=1}^{N_{ij}}y_{ijk}-rN_{ij}\tilde{\alpha}^{init}_{i}-\tilde{\sigma}^{init}_{\epsilon}\sum\limits_{j=1}^{r}\sum\limits_{k=1}^{N_{ij}}z_{ijk}^{\epsilon}}{\sum\limits_{j=1}^{r}\sum\limits_{k=1}^{N_{ij}}x_{j}\exp(\hat{\sigma}_{\eta}z_{ijk}^{\eta})},\quad i=1,...,q.$ (3)

### v3

@@ 1 Introduction
A critically important problem in interlaboratory calibration is the accurate interval estimation of the unknown true concentration of an analyte in the context of quantitative analysis in analytical chemistry. The problem is especially challenging when dealing with heterocedastic measurement errors at different concentration levels of the analyte. In practice, analytical measurements are routinely used for various purposes, often without considering the associated level of uncertainty. For instance, in environmental monitoring programs, the concentration of low-level environmental pollutants in a single sample is directly compared with regulatory standards without accounting for the uncertainty in those measurements (see Gibbons and Coleman (1994) for reviews of statistical issues in environmental monitoring applications). The uncertainty can be quantified by providing confidence bounds for the true concentration level, given measured concentrations obtained under a hierarchical structure, such as concentrations measured across different laboratories. It is common in environmental monitoring to use multiple laboratories for routine monitoring or to cross-validate an anomalous result. This nested structure, where the laboratory is considered a random factor, complicates the computation of confidence bounds for the calibration problem.
@@
Previous work in this area for linear models includes studies by Osborne (1991); Eberhardt and Mee (1994); Johnson and Krishnamoorthy (1996); Cameron et al. (2022). In addition, data analysis using the Rocke-Lorenzato model and its derivatives typically employs likelihood-based methods, maximizing the integrated likelihood to estimate parameters and their asymptotic variances. Bhaumik and Gibbons (2005) obtained moment-based estimators of the model parameters, studied their properties, and used them to construct confidence intervals using data from multiple laboratories. However, conventional large-sample confidence intervals derived from these methods often exhibit inadequate coverage probabilities, unless sample sizes are large. In response to this limitation, Zhao et al. (2021) proposed a modified likelihood ratio approach using a parametric bootstrap based on a standardized signed log-likelihood ratio test to improve confidence interval accuracy. In spite its improved performance, the coverage probabilities were still unsatisfactory for some parametric choices when multiple laboratories are involved. We shall thus explore a fiducial approach under the Rocke-Lorenzato model and its generalizations based on measurements made by a single laboratory or split samples from multiple laboratories. Based on fairly extensive simulations, we demonstrate that the fiducial approach outperforms all other available approaches, including the modified likelihood ratio procedure due to Zhao et al. (2021) and bootstrap-based approaches, in terms of meeting the coverage probability requirement, and yet providing confidence intervals having comparable or better expected widths. We also show that the proposed fiducial approach not only has better accuracy, but is also computationally much more efficient than some of the competing methods, especially the methodology developed in Zhao et al. (2021). The fiducial approach has actually seen renewed interest in recent years; Hannig et al. (2016) and Bhaumik et al. (2021). Earlier, Weerahandi (1993), Xu and Wang (2008), Krishnamoorthy and Mathew (2003) and Bhaumik et al. (2013) developed and successfully applied the fiducial approach, referring to a generalized pivotal quantity. We note that the Rocke-Lorenzato model and its derivatives have random effects that enter the model in a non-linear fashion, and we have been able to develop the fiducial approach in this non-linear scenario. We have also explored the performance of the point estimate by the fiducial mode.
@@ 2 Motivating Examples
[…]
The datasets are presented as Tables F.5 and F.6 in the Section 6 of the supplementary material.
@@ 4 Fiducial Approach
The notion of fiducial inference is based on a fiducial quantity, introduced by Fisher in the 1930s. Unlike Bayesian inference, it provides probabilistic statements about the unknown parameters without relying on prior distributions. The fiducial pivotal quantity is a function of random variables and corresponding observed values such that (i) the distribution of the fiducial pivotal quantity given the data does not depend on any unknown parameters, and (ii) the observed value of the fiducial pivotal quantity obtained by replacing each random variable by the corresponding observed data is free of the nuisance parameters, and is often equal to the parameter of interest. Applications based on fiducial quantities are available in numerous articles. In particular, in the context of lognormal distributions, such applications are given in Krishnamoorthy and Mathew (2003), Bebu and Mathew (2008), and Bhaumik et al. (2021).
@@
Remark: Interest in the fiducial approach was revived by the seminal works of Tsui and Weerahandi (1989) and Weerahandi (1993); a fiducial quantity was referred to as a generalized pivotal quantity in the latter work. Several later authors have developed and successfully applied the fiducial approach, referring to the fiducial quantity as a generalized pivotal quantity; see, for example, the articles by Krishnamoorthy and Mathew (2003), Bebu and Mathew (2008), Xu and Wang (2008), Bhaumik et al. (2013) and the book by Krishnamoorthy and Mathew (2009). The fiducial pivotal quantities we have considered in our work are in the spirit of the generalized pivotal quantities considered by these authors. We have referred to them as fiducial quantities. However, what we have developed are approximate fiducial quantities since their implementation relies on the consistency of the maximum likelihood estimate of the model parameter $\sigma_{\eta}$. In what follows, we derive fiducial quantities for each model parameter and for an unknown true concentration.
4.1 Choice of Structural Equations to Find Fiducial Pivotal Quantities and Non-Uniqueness.
We acknowledge that, in general, fiducial inference is not invariant to the particular choice of structural equations used to construct fiducial pivotal quantities, and thus different valid constructions can lead to different fiducial distributions. A more canonical choice of structural equations is available only in special settings, most notably linear mixed-effects models under linearity and balanced-data assumptions, where the construction is guided by low-dimensional sufficient statistics and yields standard generalized confidence interval (GCI) forms (see Krishnamoorthy and Mathew (2009)). Once data becomes unbalanced, even within the linear mixed-model family, such sufficient-statistic reductions are typically unavailable, and it becomes difficult to define a unique “best” structural equation or to establish optimality properties for a particular choice. Our problem inherits this challenge because we work within a nonlinear framework and encounter unbalanced data structure across laboratories and concentrations, so no universally preferred fiducial construction becomes available. Nevertheless, our specification is not arbitrary: it is chosen to align each parameter with the empirical feature it governs and to preserve computational scalability. In what follows, we will see that for fixed $(\sigma_{\eta}^{2},\sigma_{\epsilon}^{2})$, lab-specific parameters $(\alpha_{i},\beta_{i})$ have linear relations with the measurements $y_{ijk}$s and transformed measurement $y_{ijk}/x_{j}$s ($x_{j}\neq 0$) within lab $i$, which motivates us to provide structural equations based on within-lab averages leading to closed-form solutions for $\alpha_{i}$ and $\beta_{i}$. In contrast, the variance components $(\sigma_{\eta}^{2},\sigma_{\epsilon}^{2})$ primarily control (i) within-lab variability across concentrations and (ii) within-concentration variability across labs; hence, we provide solutions using aggregated within-lab and within-concentration sums of squares. This choice yields an interpretable and valid fiducial construction while requiring numerical solution only for the two variance parameters, making the procedure fast and scalable as the number of labs and concentrations increases.
@@ 4.2 Fiducial Quantities for Model Parameters
Computational motivation and initialization.
In principle, once a set of structural equations is specified, fiducial quantities can be obtained by jointly solving the full system of structural equations. In our setting, the structural equations in the following Sections 4.2.2-4.2.4 form a coupled nonlinear system with $(2q+2)$ unknowns (all $\{\alpha_{i}\}$, $\{\beta_{i}\}$, $\sigma_{\eta}^{2}$, and $\sigma_{\epsilon}^{2}$). Solving such a high-dimensional nonlinear system numerically at each fiducial draw can be computationally expensive and can scale poorly with the number of laboratories and concentrations. To address this, we adopt a computationally efficient construction that remains consistent with the fiducial definition by forming structural equations in blocks $\{\alpha_{i}\}$, $\{\beta_{i}\}$, and $\{\sigma_{\eta}^{2},\sigma_{\epsilon}^{2}\}$ and, for each block, plugging in currently available fiducial quantities for the remaining parameters. This block-wise strategy naturally requires an initialization, because at the outset, fiducial quantities for the other blocks are not yet available when constructing fiducial pivotal quantities for $\{\alpha_{i}\}$ and $\sigma_{\epsilon}^{2}$. We therefore use the measurements at zero concentration to obtain initial fiducial pivotal quantities for $\alpha_{i}$ and $\sigma_{\epsilon}^{2}$, following the same motivation as Bhaumik and Gibbons (2005). We use low concentration observations (i.e. $x=0$) to form method-of-moments type estimates of these quantities. These initial fiducial pivotal quantities are then used as inputs to the subsequent structural equations for $\{\beta_{i}\}$ and $(\sigma_{\eta}^{2},\sigma_{\epsilon}^{2})$, and finally $\{\alpha_{i}\}$ and $\sigma_{\epsilon}$ are updated using the full-data structural equations (equation (5) and equations (6)-(7) respectively). We emphasize that this two-stage step is not an approximation; rather, it is a valid fiducial construction that enables scalable computation while allowing $\alpha_{i}$ and $\sigma_{\epsilon}$ to borrow information from the entire dataset (including the nonzero concentrations), which can improve the stability of their inference and the downstream inference for unknown concentrations.
@@ 4.2.1 Initial Fiducial Quantities for $\alpha_{i}$ and $\sigma_{\epsilon}$
First, we obtain the fiducial quantities of $\alpha_{i}$, i=1,,qi=1,...,q, and $\sigma_{\epsilon}$ based on the data corresponding to $x_{j}=0$. In the next subsections, we update these fiducial quantities based on the complete data (corresponding to all concentrations $x_{j}$) to obtain updated fiducial quantities of $\alpha_{i}$, i=1,,qi=1,...,q, and $\sigma_{\epsilon}$. We assume $x_{1}=0$ and for $j=1$ the model (2) can be written as $y_{i1k}=\alpha_{i}+\sigma_{\epsilon}z_{i1k}^{\epsilon}.$ By taking an average over all replications from laboratory $i$ we obtain, $\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}=\alpha_{i}+\sigma_{\epsilon}\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}.$ Thus, we obtain, $\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}=\sigma^{2}_{\epsilon}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}.$ By summing both sides over all laboratories, so as to utilize information across all labs, we get $\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}=\sigma^{2}_{\epsilon}\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}$. Thus, an initial fiducial pivotal quantity for $\sigma_{\epsilon}$, say $\tilde{\sigma}^{init}_{\epsilon}$, can be obtained by solving the above equation for $\sigma_{\epsilon}$: $\tilde{\sigma}^{init}_{\epsilon}=\sqrt{\frac{\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}}{\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}}}.$ In the expression for $\tilde{\sigma}^{init}_{\epsilon}$ given above, the $y_{i1k}$s are to be treated as fixed, and the random quantities are the random variable $z_{i1k}^{\epsilon}$. An initial fiducial pivotal quantity for $\alpha_{i}$ can be obtained as $\tilde{\alpha}^{init}_{i}=\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}-\tilde{\sigma}^{init}_{\epsilon}\times\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}.$
Remark. In the expression for $\tilde{\sigma}^{init}_{\epsilon}$ given above, the quantity $\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(z_{i1k}^{\epsilon}-\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}z_{i1k}^{\epsilon}\right)^{2}$ clearly has a chisquare distribution. However, while implementing the fiducial methodology, we generate the independent standard normal quantities $z_{i1k}^{\epsilon}$. The reason is that these quantities also appear in later expressions; see for example equation (3).
@@ 4.2.2 Fiducial Quantities for $\beta_{i}$
We construct a total of $q$ structural equations, one for each $\beta_{i}$, i=1,,qi=1,...,q. These equations are linear in each $\beta_{i}$, i=1,,qi=1,...,q; thus, solutions are obtained in closed forms. We use (2) to construct structural equations for each $\beta_{i}$, i=1,,qi=1,...,q, by averaging (2) over all replications corresponding to different concentrations, $\frac{1}{r}\sum\limits_{j=1}^{r}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}y_{ijk}=\alpha_{i}+\beta_{i}\frac{1}{r}\sum\limits_{j=1}^{r}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}x_{j}\exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\frac{1}{r}\sum\limits_{j=1}^{r}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}z_{ijk}^{\epsilon},\ i=1,...,q.$ Parameters $\alpha_{i}$, i=1,,qi=1,...,q, and $\sigma^{2}_{\epsilon}$ in these structural equations are replaced by their initial fiducial quantities, $\tilde{\alpha}^{init}_{i}$, i=1,,qi=1,...,q, and $\tilde{\sigma}^{init}_{\epsilon}$ respectively. We replace $\sigma^{2}_{\eta}$ with its consistent point estimate $\hat{\sigma}^{2}_{\eta}$ which can be obtained by the maximum likelihood or the method of moments proposed by Bhaumik and Gibbons (2005). Using $\hat{\sigma}^{2}_{\eta}$, the fiducial quantities that we exhibit for the $\beta_{i}$, i=1,,qi=1,...,q, are only approximate fiducial quantities; however, we shall henceforth refer to the approximate fiducial quantities simply as fiducial quantities. Thus, the fiducial pivotal quantities for $\beta_{i}$, i=1,,qi=1,...,q, can be obtained as
$\tilde{\beta}_{i}=\frac{\sum\limits_{j=1}^{r}\sum\limits_{k=1}^{N_{ij}}y_{ijk}-\sum\limits_{j=1}^{r}N_{ij}\tilde{\alpha}^{init}_{i}-\tilde{\sigma}^{init}_{\epsilon}\sum\limits_{j=1}^{r}\sum\limits_{k=1}^{N_{ij}}z_{ijk}^{\epsilon}}{\sum\limits_{j=1}^{r}\sum\limits_{k=1}^{N_{ij}}x_{j}\exp(\hat{\sigma}_{\eta}z_{ijk}^{\eta})},\quad i=1,...,q.$ (3)